OpenStudy (anonymous):

Determine the odds against the spinner landing on green? 3-red, 2-yellow, 2-blue, 1-green

OpenStudy (anonymous):

Did u get that...

OpenStudy (anonymous):

There are 8 possible things that it could land on. Only 1 of those is green, that means that the chance of landing on green is 1/8. How do you find the chance of it not landing on green if you know what the probability of it landing on green is?

OpenStudy (anonymous):

Does this determine the odds against it landing green..

OpenStudy (anonymous):

I think you are getting this, but it's unclear from your last question. If you know the chance that it IS green, then what is the chance that it is NOT green?

OpenStudy (anonymous):

1 to 7 or 7 to 1

OpenStudy (anonymous):

there are 1 out of 8 ways that it can be green, so there are 7 out of 8 ways it will be "not green" Do you have to give your answer as "odds"? Or normal probability?

OpenStudy (anonymous):

2le gave you the probability answer, but that's not quite the same as "odds".

OpenStudy (anonymous):

odds

OpenStudy (anonymous):

I think you're right then... 7 to 1 that it will be "not green"... there are 7 ways to get "not green" compared to 1 way to be "green"

OpenStudy (anonymous):

but I'm not totally sure if it's 7 to 1 or 1 to 7

OpenStudy (anonymous):

in fact, I'm sure it isn't 7 to 1... that means the opposite... a low chance of being not green

OpenStudy (anonymous):

sorry, confused myself!! :)

OpenStudy (anonymous):

Now would this be the same concept? Determine the probability that the spinner lands on yellow? 3-red, 2-yellow, 2-blue, 1-green

OpenStudy (anonymous):

maybe I confused myself twice. just looked this up and it said ODDS = "chances for" : "chances against" which would seem like the odds for getting "not green" might be 7:1

OpenStudy (anonymous):

I am better at probability :) probability of yellow = # of ways to get yellow / total # of outcomes = 2 / 8 =1 / 4 or 25%

OpenStudy (anonymous):

Here's something to check on odds, now that I have confused things... http://www.math-magic.com/probability/prob_to_odds.htm

OpenStudy (anonymous):

Tkx u!!!

OpenStudy (anonymous):

you're welcome :)